A 210 kg mass initially at rest on a smooth horizontal plane is acted upon by a force F parallel with the plane. If F is 1.4 tN where t is time in seconds, find for an interval of 30 seconds from rest: (a) the impulse (b) the velocity (c) the average force producing the same impulse.
1
Expert's answer
2020-02-13T09:04:05-0500

Impulse is given by "\intop Fdt"

(a) Impulse= "\intop_0^{30}" "1.4tdt" "=1.4\times450=630Ns"


(b) Impulse= change in momentum "=\Delta p=m(v_f-v_i)"

"630=210(v-0) \newline v=3m/sec"


(c) "Impulse = avg. force \times time"

"630=F_{avg}\times30 \newline
F_{avg}=21 N"